OpenStudy (anonymous):

find the value of x>1 for which the function F(x)=integral x to x^2 [(1/t)*{ln([1-t]/32}]dt is increasing and decreasing

OpenStudy (anonymous):

find the value of x>1 for which the function \[F(x)= \int\limits_{x}^{x^2}(1/t)*\ln [(1-t)/32]dt \] is increasing and decreasing

OpenStudy (anonymous):

F'(x) = 2x [ (1/x^2) ln(1-x^2)/32 ] find where f'(x) is positive or negative

OpenStudy (anonymous):

@sourwing i think there should be an \(x\) in the denominator, not that it makes much difference

OpenStudy (anonymous):

ah yes, I was just noticed the x lol

OpenStudy (anonymous):

i am also getting a slightly different derivative, but my algebra could be off

OpenStudy (anonymous):

let me do the derivative again

OpenStudy (anonymous):

i get as a first step \[\frac{\log(1-x^2)}{16x}-\frac{\log(1-x)}{32x}\]

OpenStudy (anonymous):

as a second step \[\log((1-x^2)^2)-\log(1-x)\] ignoring the denominator which is positive in any case

OpenStudy (anonymous):

but like i said, my algebra could be not so good at this hour

OpenStudy (anonymous):

F(x) = ∫{x->2} (1/t) ln(1-t^2)/32 + ∫[2->x^2] (1/t) ln(1-t^2)/32 F(x) = -∫{2->x} (1/t) ln(1-t^2)/32 + ∫[2->x^2] (1/t) ln(1-t^2)/32 F'(x) = - (1/x) ln(1-x)/32 + (1/x^2) ln(1-x^2)/32 * 2x

OpenStudy (anonymous):

no matter, i think we lost the person answering the question

OpenStudy (anonymous):

yeah that is what i got too then cancelled an \(x\) in the second term, added up the fractions and got the numerator i wrote above

OpenStudy (loser66):

@satellite73 is it not that (integral (f (x)))' = f(x)?

OpenStudy (anonymous):

@loser you have an \(x\) at both ends of the limit of integration

OpenStudy (anonymous):

\[\int _a^{x^2}f(t)dt)'=f(x^2)\times 2x\] by the chain rule

OpenStudy (anonymous):

F'(x) = ln(1-x)/(32x)

OpenStudy (anonymous):

oh wait no

OpenStudy (anonymous):

i get this as a first step \[\frac{\log(1-x^2)}{16x}-\frac{\log(1-x)}{32x}\]

OpenStudy (anonymous):

by \(\frac{2x}{32x^2}=\frac{1}{16x}\)

OpenStudy (anonymous):

yep got the samething

OpenStudy (anonymous):

@Loser66 the derivative of \[\int_a^xf(t)dt\] is \(f(x)\) but via the chain rule the derivative of \[\int_a^{g(x)}f(t)dt\] is \[f(g(x))g'(x)\]

OpenStudy (anonymous):

then adding up the fractions gave me \[\log((1-x^2)^2)-\log(1-x)\] as a numerator

OpenStudy (anonymous):

property of logs gives \[\log((1+x)^2(1-x))\]

OpenStudy (anonymous):

yw @Loser66 also you have an \(x\) as the lower limit of integration, so you have to subtract off \(f(x)\)

OpenStudy (anonymous):

@satellite73 shoult it be ln[ (1-x^2)^2 / (1-x) ]?

OpenStudy (anonymous):

it should be but \[\frac{1-x^2}{1-x}=1+x\] so i cancelled one of them

OpenStudy (anonymous):

ahh ..

OpenStudy (anonymous):

not really sure it makes it easier, but i figured it probably would question is still for which \(x\) is that thing positive

OpenStudy (anonymous):

i guess all the work is solving \[(1+x)^2(1-x)>1\] for \(x\) but maybe there is a better method

OpenStudy (anonymous):

oh no, it works out well get \[-x^3-x^2+x > 0\]

OpenStudy (anonymous):

end up solving a quadratic equation all this work and the person is gone too bad

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

don't we need to take the domain into account?

OpenStudy (anonymous):

the denominator is unimportant since \(x>1\) is given

OpenStudy (anonymous):

i think you end up with the golden mean for this one

OpenStudy (anonymous):

\[x (x^2+x-1)<0\] is a first step

OpenStudy (anonymous):

well, I thought F(x) is only differentiable on (-1,0) U (0,1)?

OpenStudy (anonymous):

not sure why problem states \(x>1\) so in integrand is defined for all \(x\) if \(x>1\)

OpenStudy (anonymous):

I'm pretty sure there is a typo.

OpenStudy (anonymous):

well, i am not going to fret about it since i am just fooling around but i did think it was rather cool that you end up with the golden mean \[1<x<\frac{-1+\sqrt5}{2}\]

OpenStudy (anonymous):

urg... screw it :/ i'm off playing game lol

OpenStudy (anonymous):

where did u got \[[\ln (1-x ^{2})/16x]-[\ln (1-x)/32x]\] from pl explain

OpenStudy (anonymous):

pl help me how u found f'(x)